OpenStudy (anonymous):

Use the intercepts to graph the equations x+5y=10

OpenStudy (saifoo.khan):

When y intercpts put x = 0 When x intercpts put y = 0

OpenStudy (anonymous):

what are you saying saifoo

OpenStudy (anonymous):

you saying there is no answer or graph 0 y and 0 x

OpenStudy (saifoo.khan):

Let me explain, just a sec.

OpenStudy (saifoo.khan):

When y intercpts put x = 0 x+5y=10 0 +5y = 10 5y = 10 y = 2 (0,2) <------ y intercept. When x intercpts put y = 0 x+5y=10 x + 5(0) = 10 x + 0 = 10 x = 10 So, (10,0) <----------- x intercept.

OpenStudy (anonymous):

oh ok i see

OpenStudy (saifoo.khan):

That's the easiest method which you will never ever forget. :D